# An Information Theoretic Perspective on Conformal Prediction

### v1

@@ 5 Side Information
With the information theoretical interpretation of conformal prediction, we can translate various intuitions from information theory, for example about different type of channels or network information theory, to conformal prediction. In this section, we consider the notion of side information.
Let $X$ be the input covariates, $Y$ be the target variable, $Z$ be some side information about the task and let $Q_{Y|X}$ be the model which we use to perform conformal prediction. As we can relate CP with $Q_{Y|X}$ to an upper bound on the conditional entropy $H(Y|X)$, we would like to do the same for the case of the conditional entropy when side information is available, i.e., $H(Y|X,Z)$. Since we know that the conditional entropy directly affects the expected set size, i.e., the inefficiency of CP, and given that $H(Y|X)\geq H(Y|X,Z)$ we can expect that with the additional side information the inefficiency of CP will decrease. We can take side information into account by defining conformity scores as a function of $Q_{Y|X,Z}$ instead of $Q_{Y|X}$. A simple way to do that would be via the Bayes rule
@@ 5.1 The Distributed Learning Setting
Consider the case where we have a dataset that is distributed among a set of $m$ devices and we want to run conformal training for a global model $Q_{Y|X}$. Further, assume it is difficult to gather the data at a central location (e.g., due to privacy reasons), and thus we have to work with the data staying locally on each device. An example of this would be Federated Learning (or FL) (McMahan et al., 2017). In this case, if $Z\in\{1,\ldots,m\}$ identifies the device, the entropy $H(Y|X)$ can be expressed as
@@
where the first inequality is due to $Z$ being discrete and having non-negative entropy and the second is due to Gibbs inequality with $Q_{Z|X}$ being an auxiliary model trained to predict the user id $Z=z$ given input $X=x$. A similar upper bound has been considered before in a federated setting (Louizos et al., 2024). With this upper bound we get
@@
This gives us an upper bound on the entropy of the entire population that decomposes into a sum of local functions, one for each client, only requiring local information. As a result, we can easily perform conformal training for $Q_{Y|X}$ by optimizing this upper bound in the federated setting with, e.g, federated averaging (McMahan et al., 2017). At test time, we can consider the device ID $z$ as side information for CP. To this end, we can either train in parallel a model $Q_{Z|X,Y}$ and use 8 with the global model $Q_{Y|X}$ at test time in order to compute an adjusted model $Q_{Y|X,Z}$ or obtain $Q_{Y|X,Z}$ by fine tuning the global model $Q_{Y|X}$ with local data.
@@ 6 Related Work
Conformal prediction, a powerful framework for uncertainty estimation developed by Vovk and collaborators (Vovk et al., 2005; Shafer & Vovk, 2008) has recently witnessed a wide adoption in many fields, e.g., healthcare (Papadopoulos et al., 2009; Alnemer et al., 2016; Lu et al., 2022b, a) and finance (Wisniewski et al., 2020; Bastos, 2024). The marriage of conformal prediction and machine learning has been especially fruitful, and since the seminal work by Vovk et al. (2005), many extensions and applications of CP to ML have been proposed in the literature. Notably, concerning topics such as survival analysis (Candès et al., 2021), treatment effect evaluation (Lei & Candès, 2021), classification (Gupta et al., 2020; Angelopoulos et al., 2020) and regression (Romano et al., 2019) settings, risk-controlling (Angelopoulos et al., 2021; Bates et al., 2021), and covariate shift (Tibshirani et al., 2019), to mention some.
The original conformal prediction framework, often referred to as full or transductive conformal prediction (Vovk et al., 2005), requires retraining the predictive function for each new test data point and each of its possible labels. This quickly becomes too computationally intensive for modern machine learning methods, which makes the split conformal prediction variant (Papadopoulos et al., 2002; Lei et al., 2015) particularly appealing in practice. SCP allows us to transform any pointwise predictor, e.g., a large pretrained ML model, into a set predictor encoding a rigorous notion of uncertainty in the size of each prediction set (Angelopoulos & Bates, 2021; Angelopoulos et al., 2021; Bates et al., 2021). Unfortunately, this comes at the cost of allocating a chunk of the data set solely for calibration, which reduces the amount of data available for training model.
@@
To the best of our knowledge, our work represents the first attempt to bridge CP and information theory. Among other things, this allows us to build on the conformal training ideas of Bellotti (2021) and Stutz et al. (2022), deriving principled learning objectives that generalize their approaches, dispense with some their hyperparameters and result in more efficient prediction sets. Moreover, we also empirically demonstrate that our conformal training objectives provide a strong enough learning signal to train complex architectures from scratch, with strong results on ResNet-34 and ResNet-50 (He et al., 2016) fitted on CIFAR10 and CIFAR100, respectively. In contrast, the previous state-of-the-art approach, ConfTr (Stutz et al., 2022), struggles in those settings (see experiments in Section 7) and required pretrained models for consistent results (Stutz et al., 2022). Further, our information theoretical interpretation of CP provides a new simple and effective mechanism to leverage side information in split conformal prediction. We are unaware of any other approaches to treat side information inside the conformal prediction framework in the literature.
@@
On the information theory side, the notion of $f$-divergences and related inequalities have been discussed in many different works. The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s, as in, for instance, (Ali & Silvey, 1966; Csiszár, 1967; Morimoto, 1963). A key $f$-divergence inequality is the data processing inequality—see (Sason & Verdú, 2016; Sason, 2019) for an extensive survey. It provides a unified way of obtaining many classical and new results, for example Fano’s inequality (Fano, 1949). The tightness of the data processing inequalities is discussed in terms of Bregman’s divergence in (Collet, 2019; Liese & Vajda, 2006) and in terms of $\chi^{2}$-divergence in (Sason, 2019). List decoding, which is closely connected to CP, was introduced in context communication design by Elias (Elias, 1957) and Wozencraft (Wozencraft, 1958). A generalization of Fano’s inequality to list decoding was given in (Csiszár & Körner, 2011) in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in (Raginsky & Sason, 2013). Variable-size list decoding was discussed in (Sason, 2019), using ideas first introduced in (Polyanskiy et al., 2010) and (Liu et al., 2017). A selection of relevant inequalities for list decoding can be found in (Sason, 2019).
@@
Table 1: Inefficiency with both THR and APS methods for each objective and data set for a target $\alpha=0.01$. We report mean and standard deviation across 10 different calib./test splits and show in bold all values within one standard deviation of the best result. Lower is better.
@@
CE $\mathbf{2.15_{\pm 0.21}}$ $2.41_{\pm 0.12}$ $2.18_{\pm 0.15}$ $2.24_{\pm 0.21}$ $\mathbf{2.01_{\pm 0.06}}$ $3.27_{\pm 0.16}$ $\mathbf{1.81_{\pm 0.20}}$ $2.41_{\pm 0.25}$ $21.11_{\pm 1.17}$ $28.11_{\pm 2.38}$
ConfTr $\mathbf{2.10_{\pm 0.16}}$ $\mathbf{2.10_{\pm 0.16}}$ $\mathbf{1.69_{\pm 0.06}}$ $1.87_{\pm 0.08}$ $\mathbf{1.95_{\pm 0.05}}$ $\mathbf{2.32_{\pm 0.06}}$ $9.77_{\pm 0.02}$ $9.98_{\pm 0.00}$ $41.87_{\pm 1.71}$ $29.57_{\pm 1.09}$
@@
Fano $\mathbf{2.08_{\pm 0.21}}$ $\mathbf{2.08_{\pm 0.13}}$ $\mathbf{1.73_{\pm 0.03}}$ $1.87_{\pm 0.06}$ $\mathbf{2.02_{\pm 0.08}}$ $2.68_{\pm 0.10}$ $2.00_{\pm 0.10}$ $2.35_{\pm 0.10}$ $34.51_{\pm 2.33}$ $35.11_{\pm 1.53}$
MB Fano $\mathbf{2.15_{\pm 0.11}}$ $\mathbf{2.37_{\pm 0.24}}$ $\mathbf{1.77_{\pm 0.06}}$ $2.25_{\pm 0.12}$ $2.86_{\pm 0.22}$ $3.67_{\pm 0.15}$ $\mathbf{1.70_{\pm 0.04}}$ $\mathbf{2.01_{\pm 0.08}}$ $\mathbf{17.43_{\pm 0.93}}$ $21.27_{\pm 1.38}$
DPI $\mathbf{2.18_{\pm 0.13}}$ $2.52_{\pm 0.12}$ $\mathbf{1.75_{\pm 0.08}}$ $2.06_{\pm 0.08}$ $\mathbf{1.98_{\pm 0.07}}$ $4.14_{\pm 0.12}$ $\mathbf{1.71_{\pm 0.07}}$ $\mathbf{1.99_{\pm 0.10}}$ $\mathbf{16.03_{\pm 1.02}}$ $\mathbf{16.86_{\pm 1.19}}$
7 Experiments
In this section, we empirically study two applications of our theoretical results, namely conformal prediction with side information and conformal training with our upper bounds on the conditional entropy as optimization objectives. We focus our experiments on classification tasks, since this is the most common setting in previous works in conformal training (Bellotti, 2021; Stutz et al., 2022).
7.1 Conformal Training
[…]
For each data set, we use the default train and test splits but transfer 10% of the training data to the test data set. We train the classifiers only on the remaining 90% of the training data and, at test time, run SCP with 10 different calibration/test splits by randomly splitting the enlarged test data set. Further, for each data set, objective and SCP method, we run grid search to find the best values for batch size, learning rate, temperature (smoothing parameter in the assignment of labels to prediction sets) and steepness (smoothing parameter in diffsort (Petersen et al., 2022), the differentiable sorting operator we used). For the remaining ConfTr hyperparameters, we used the best values reported in (Stutz et al., 2022). To avoid overfitting to the test data, the grid search was done solely on the 90% of the training data not used for testing. See Appendix F for a detailed account of the experimental settings, including additional results, model architectures, and optimal hyperparameters.
[…]
Table 2: Inefficiency on CIFAR100 for different $\alpha$ targets at test time, averaged across 10 random calib./test splits. All methods optimized only for $\alpha{=}0.01$. Lower is better.
Method $\alpha=0.01$ $\alpha=0.05$ $\alpha=0.1$
THR APS THR APS THR APS
CE $21.11_{\pm 1.17}$ $27.41_{\pm 2.16}$ $6.40_{\pm 0.37}$ $9.95_{\pm 0.56}$ $3.24_{\pm 0.14}$ $4.97_{\pm 0.20}$
ConfTr $41.87_{\pm 1.71}$ $29.57_{\pm 1.09}$ $12.06_{\pm 0.27}$ $21.22_{\pm 0.82}$ $7.15_{\pm 0.26}$ $14.33_{\pm 0.70}$
[…]
Fano $34.51_{\pm 2.33}$ $35.11_{\pm 1.53}$ $15.14_{\pm 0.56}$ $18.30_{\pm 0.42}$ $8.84_{\pm 0.33}$ $10.86_{\pm 0.23}$
MB Fano $\mathbf{17.43_{\pm 0.93}}$ $21.27_{\pm 1.38}$ $5.95_{\pm 0.35}$ $9.33_{\pm 0.27}$ $3.18_{\pm 0.12}$ $5.57_{\pm 0.2}$
DPI $\mathbf{16.03_{\pm 1.02}}$ $\mathbf{16.86_{\pm 1.19}}$ $\mathbf{5.21_{\pm 0.17}}$ $\mathbf{7.06_{\pm 0.24}}$ $\mathbf{2.71_{\pm 0.11}}$ $\mathbf{4.14_{\pm 0.12}}$
In all experiments, we optimize the conformal training objectives for a target coverage rate of 99%, i.e., $\alpha=0.01$. It is then important to evaluate whether the performance of the resulting models deteriorate at different coverage rates, “overfitting” to the value of $\alpha$ used for training. In Table 2, we see that is not the case, since our bounds still outperform the baselines—including the cross-entropy loss, which is agnostic to the desired coverage rate—at different $\alpha$ values. Table 2 also confirms the observation in the original paper (Stutz et al., 2022) that the performance of ConfTr deteriorates for higher $\alpha$ values. For $\alpha=0.1$, both ConfTr methods produce prediction sets that are more than two times larger than those obtained via the CE loss.
@@ 7.2 Side Information
As a first experiment, we consider datasets for which there exists a natural grouping of the labels and use the group assignment as side information $Z$. In CIFAR100, there is a disjoint partition of the 100 classes into 20 superclasses, so we define $z$ as the superclass to which example $(x,y)$ belongs. In EMNIST, $z$ indicates whether the example is a digit, uppercase or lowercase letter. We train an auxiliary model $R_{Z|X,Y}$ and, at test time, assume access to side information $z$ to recompute class probabilities $Q_{Y|X,Z=z}$ from the original classifier $Q_{Y|X}$ as in equation (8).
We report results in Table 3, where we consider two different scenarios. The first is the standard SCP setting, where we evaluate the inefficiency of THR and APS methods, with side information $Z$ observed for 10, 30 and 100% of the instances. In this case, we redefine the classifier $f$ as in (9) to account for when $Z$ is missing, but otherwise the SCP process remain unchanged. The second scenario is Mondrian or Group-Balanced CP (Vovk et al., 2005), where one splits ${\mathcal{D}}_{cal}$ into different groups and runs CP for each of them individually. In this setting, we group the calibration data points according to $Z$ and base the score function on $Q_{Y|X,Z}$. In all cases, taking the side information into account reduced the inefficiency considerably.
Table 3: Inefficiency for each method and dataset for a target $\alpha=0.01$ with side information. The side information is the superclass assignment for CIFAR100 and whether the class is a digit / uppercase letter / lowercase letter for EMNIST.
@@ 7.3 Federated Learning (FL)
Table 4: Inefficiency of conformal prediction in the federated setting with THR and APS for each dataset and for a target $\alpha=0.01$. We report mean and standard deviation of the inefficiency of the global federated model, with and without side information, across 10 different calib./test splits and show in bold all values within one standard deviation of the best result. Lower is better.
@@
THR THR${}_{+\text{side}}$ APS APS${}_{+\text{side}}$ THR THR${}_{+\text{side}}$ APS APS${}_{+\text{side}}$ THR THR${}_{+\text{side}}$ APS APS${}_{+\text{side}}$
CE $2.91_{\pm 0.02}$ $2.46_{\pm 0.02}$ $3.69_{\pm 0.03}$ $3.14_{\pm 0.04}$ $2.73_{\pm 0.04}$ $2.30_{\pm 0.06}$ $\mathbf{2.83_{\pm 0.07}}$ $2.43_{\pm 0.06}$ $55.41_{\pm 1.09}$ $52.31_{\pm 1.03}$ $64.73_{\pm 0.34}$ $62.67_{\pm 3.68}$
ConfTr $4.60_{\pm 0.05}$ $3.30_{\pm 0.02}$ $6.14_{\pm 0.04}$ $5.25_{\pm 0.04}$ $10.00_{\pm 0.00}$ $10.00_{\pm 0.00}$ $10.00_{\pm 0.00}$ $10.00_{\pm 0.00}$ $\mathbf{45.60_{\pm 1.30}}$ $\mathbf{41.18_{\pm 1.16}}$ $55.18_{\pm 2.10}$ $47.58_{\pm 1.48}$
@@
Fano $\mathbf{2.63_{\pm 0.02}}$ $2.37_{\pm 0.02}$ $3.12_{\pm 0.04}$ $2.72_{\pm 0.03}$ $\mathbf{2.39_{\pm 0.07}}$ $\mathbf{2.07_{\pm 0.07}}$ $\mathbf{2.73_{\pm 0.07}}$ $\mathbf{2.39_{\pm 0.06}}$ $\mathbf{47.91_{\pm 1.20}}$ $\mathbf{41.19_{\pm 1.02}}$ $\mathbf{46.95_{\pm 0.67}}$ $\mathbf{42.75_{\pm 0.91}}$
MB Fano $2.84_{\pm 0.04}$ $2.25_{\pm 0.03}$ $4.75_{\pm 0.03}$ $\mathbf{2.43_{\pm 0.01}}$ $\mathbf{2.52_{\pm 0.07}}$ $\mathbf{2.04_{\pm 0.07}}$ $\mathbf{2.79_{\pm 0.13}}$ $\mathbf{2.33_{\pm 0.05}}$ $52.94_{\pm 1.40}$ $46.97_{\pm 1.30}$ $50.72_{\pm 1.77}$ $45.72_{\pm 1.38}$
DPI $\mathbf{2.60_{\pm 0.02}}$ $2.23_{\pm 0.01}$ $2.98_{\pm 0.03}$ $2.58_{\pm 0.02}$ $2.76_{\pm 0.07}$ $2.28_{\pm 0.03}$ $\mathbf{2.68_{\pm 0.15}}$ $\mathbf{2.22_{\pm 0.09}}$ $52.36_{\pm 0.95}$ $48.64_{\pm 0.70}$ $51.29_{\pm 1.07}$ $47.18_{\pm 1.27}$
After convergence, we ran SCP with the final global model assuming calibration and test data sets at the server, or equivalently that the clients share their scores with the server. This reflects the best inefficiency results we can hope for with the global model, as in practice we might need to resort to privacy preserving methods that are likely to hurt performance. See Appendix F for a discussion and extra results on other possible settings. We report inefficiency results for the global model in Table 4, where we observe similar trends to the centralized experiments in Table 1. Namely, ConfTr methods still perform well on EMNIST but struggle on CIFAR data—with the notable exception on CIFAR100, where ConfTr excelled—while our methods achieved consistent results across all data sets. This probably reflects the sensitivity of both ConfTr objectives to hyperparameters, which make them hard to use in practice, especially in FL where hyperparameter optimization is computationally intensive. Conversely, our methods seem more robust to such variations, as the hyperparameters found on the centralized setting seem to translate well to the federated case.
One marked difference between Tables 1 and 4 is that the simple Fano bound (6), which lagged behind the DPI bound and its model-based counterpart on the centralized setting, achieved the best results on the federated setting. We hypothesize this could be due to overfitting of the local optimization procedures to the individual data distribution of each device, which hurts the convergence of the global model. This effect is exacerbated on CIFAR100, where we have 500 devices, each of which ends with very few data points. The simple Fano bound is less vulnerable to such overfitting, since it relies on the main classifier to a much lesser degree.
Finally, in almost all cases our upper bounds outperformed the cross-entropy, reassuring the potential of conformal training. Moreover, the inclusion of side information reduced inefficiency in all settings, and markedly so in a few instances. This confirms the effectiveness of our side information approach in a complex and practically relevant scenario, like federated learning.

### v2

@@ 5 Side Information
With the information theoretical interpretation of conformal prediction, we can translate various intuitions from information theory, for example, about different types of channels or network information theory, to conformal prediction. In this section, we consider the notion of side information.
Let $X$ be the input covariates, $Y$ be the target variable, $Z$ be some side information about the task, and let $Q_{Y|X}$ be the model which we use to perform conformal prediction. As we can relate CP with $Q_{Y|X}$ to an upper bound on the conditional entropy $H(Y|X)$, we would like to do the same for the case of the conditional entropy when side information is available, i.e., $H(Y|X,Z)$. Since we know that the conditional entropy directly affects the expected set size, i.e., the inefficiency of CP, and given that $H(Y|X)\geq H(Y|X,Z)$ we can expect that with the additional side information the inefficiency of CP will decrease. We can take side information into account by defining conformity scores as a function of $Q_{Y|X,Z}$ instead of $Q_{Y|X}$. A simple way to do that would be via the Bayes rule
@@ 5.1 The Distributed Learning Setting
Consider the case where we have a dataset that is distributed among a set of $m$ devices and want to run conformal training to get a global model $Q_{Y|X}$ trained on all the data. Further, assume it is hard to gather the data at a central location (e.g., due to privacy reasons), and thus we have to work with the data staying locally on each device. An example of this would be Federated Learning (or FL) [40]. In this case, if $Z\in\{1,\ldots,m\}$ identifies the device, the entropy $H(Y|X)$ can be expressed as
@@
where the first inequality is due to $Z$ being discrete and having non-negative entropy and the second is due to Gibbs inequality with $Q_{Z|X}$ being an auxiliary model trained to predict the user id $Z=z$ given input $X=x$. A similar upper bound has been considered before in a federated setting [34]. With this upper bound we get
@@
This gives us an upper bound on the entropy of the entire population that decomposes into a sum of local functions, one for each client, only requiring local information. As a result, we can easily carry out conformal training for $Q_{Y|X}$ by optimizing this upper bound in the federated setting with, e.g., federated averaging [40]. At test time, we can take the device ID $z$ as side information. To this end, we can either train in parallel a model $Q_{Z|X,Y}$ and use (9) with the global model $Q_{Y|X}$ at test time in order to compute an adjusted model $Q_{Y|X,Z}$, or we can obtain $Q_{Y|X,Z}$ by fine-tuning the global model $Q_{Y|X}$ with local data.
@@ 6 Related Work
Conformal prediction, a powerful framework for uncertainty quantification developed by Vovk and collaborators [62, 58] has recently witnessed a wide adoption in many fields, e.g., healthcare [46, 3, 37, 36] and finance [65, 6]. The marriage of conformal prediction and machine learning has been especially fruitful. Since the seminal work by Vovk et al. [62], many extensions and applications have been proposed, covering topics such as survival analysis [10], treatment effect evaluation [30], classification [20, 4] and regression [52] settings, risk control [5, 7], and covariate shift [60].
@@
To our knowledge, our work represents the first attempt to bridge conformal prediction and information theory. Among other things, this allows us to build on the conformal training ideas of Bellotti [8] and Stutz et al. [59], deriving principled learning objectives that generalize their approaches, dispense with some their hyperparameters and result in more efficient prediction sets. Further, we empirically show that our conformal training objectives provide a strong enough learning signal to train complex architectures from scratch, with strong results on ResNet-34 and ResNet-50 [21] fitted on CIFAR10 and CIFAR100, respectively. In contrast, the previous state-of-the-art approach, ConfTr, struggles in those settings (see experiments in Section 7) and required pretrained models for consistent results [59]. Further, our information-theoretic interpretation of CP provides a new simple and effective mechanism to leverage side information in split conformal prediction. We are unaware of any other approaches to treat side information inside the conformal prediction framework in the literature.
@@
On the information theory side, the notion of $f$-divergence and related inequalities have appeared in many different works. The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s, as in, for instance, [2, 14, 42]. A key $f$-divergence inequality is the data processing inequality—see [56, 55] for an extensive survey. It provides a unified way of obtaining many classical and new results, including Fano’s inequality [18]. The tightness of the data processing inequalities is discussed in terms of Bregman’s divergence in [12, 32] and in terms of $\chi^{2}$-divergence in [55]. List decoding, which is closely connected to CP, was introduced in context communication design by Elias [17] and Wozencraft [66]. A generalization of Fano’s inequality to list decoding was given in [15] in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in [51]. Variable-size list decoding was discussed in [55], using ideas first introduced in [50] and [33]. A selection of relevant inequalities for list decoding can be found in [55].
7 Experiments
In this section, we empirically study two applications of our theoretical results, namely conformal prediction with side information and conformal training with our upper bounds on the conditional entropy as optimization objectives. We focus our experiments on classification tasks since this is the most common setting in previous works in conformal training [8, 59].
7.1 Conformal Training
[…]
@@ 7.1 Conformal Training
Table 1: Inefficiency results for conformal training in the centralized setting. We report the mean prediction set size ($\pm$ standard deviation) across 10 different calib./test splits for $\alpha=0.01$, showing in bold all values within one std. of the best result. Results for THR and APS correspond to different models trained with different hyperparameters (see Appendix G). Lower is better.
@@
CE $\mathbf{2.29_{\pm 0.18}}$ $2.50_{\pm 0.08}$ $2.39_{\pm 0.13}$ $2.41_{\pm 0.17}$ $\mathbf{2.06_{\pm 0.11}}$ $3.40_{\pm 0.18}$ $\mathbf{1.69_{\pm 0.11}}$ $2.34_{\pm 0.22}$ $19.70_{\pm 2.05}$ $26.02_{\pm 1.31}$
ConfTr $6.28_{\pm 0.71}$ $\mathbf{2.10_{\pm 0.07}}$ $\mathbf{1.73_{\pm 0.06}}$ $\mathbf{1.89_{\pm 0.09}}$ $\mathbf{1.99_{\pm 0.10}}$ $\mathbf{2.36_{\pm 0.11}}$ $9.90_{\pm 0.02}$ $9.98_{\pm 0.00}$ $32.80_{\pm 2.75}$ $40.58_{\pm 1.23}$
@@
Fano $\mathbf{2.09_{\pm 0.12}}$ $\mathbf{2.12_{\pm 0.08}}$ $\mathbf{1.70_{\pm 0.05}}$ $\mathbf{1.87_{\pm 0.05}}$ $\mathbf{2.10_{\pm 0.11}}$ $2.75_{\pm 0.14}$ $2.05_{\pm 0.05}$ $2.35_{\pm 0.10}$ $40.30_{\pm 1.10}$ $33.80_{\pm 0.93}$
MB Fano $\mathbf{2.24_{\pm 0.12}}$ $2.49_{\pm 0.19}$ $\mathbf{1.80_{\pm 0.08}}$ $2.25_{\pm 0.14}$ $\mathbf{2.01_{\pm 0.11}}$ $3.67_{\pm 0.13}$ $\mathbf{1.66_{\pm 0.09}}$ $\mathbf{1.89_{\pm 0.06}}$ $\mathbf{14.61_{\pm 0.84}}$ $21.68_{\pm 1.44}$
DPI $\mathbf{2.24_{\pm 0.17}}$ $2.64_{\pm 0.07}$ $\mathbf{1.73_{\pm 0.07}}$ $2.08_{\pm 0.06}$ $\mathbf{1.98_{\pm 0.09}}$ $4.07_{\pm 0.23}$ $\mathbf{1.64_{\pm 0.07}}$ $\mathbf{1.97_{\pm 0.08}}$ $17.55_{\pm 1.33}$ $\mathbf{17.41_{\pm 0.62}}$
In Table 1, we report the empirical inefficiency on test data considering two SCP methods, threshold CP with probabilities (or THR) [54] and APS [53]—see Appendix G.1.2 for results with RAPS [4]. In all cases, our upper bounds proved effective loss functions to train efficient classifiers end-to-end and from scratch. For the simpler data sets (MNIST, Fashion-MNIST and EMNIST), all conformal training methods achieved similar results, but both ConfTr methods proved less consistent. This is noticeable in the oftentimes sharp difference in performance between THR and APS, since even after fine-tuning the hyperparameters (see Appendix G) some of the models fail to converge properly. For the remaining and more challenging data sets, both ConfTr variants lagged behind, probably because they do not provide a strong enough signal to train ResNets from scratch (on CIFAR data sets, Stutz et al. [59] only used ConfTr to fine-tune pretrained models). A similar observation applies to the Fano bound, where the only signal to the classifier comes through the relaxed prediction set size in (6). Finally, we also observe that our conformal training objectives outperform the baselines even if $\alpha$ values at test time differ from the one used for training, as we show in Table 10 in Appendix G.1.1.
@@ 7.2 Side Information
As a first experiment, we consider datasets for which a natural grouping of the labels exists and use the group assignment as side information $Z$. In CIFAR100, there is a disjoint partition of the 100 classes into 20 superclasses, so we define $z$ as the superclass to which example $(x,y)$ belongs. In EMNIST, $z$ indicates whether the example is a digit, uppercase or lowercase letter. We train an auxiliary model $R_{Z|X,Y}$ and, at test time, assume access to side information $z$ to recompute class probabilities $Q_{Y|X,Z=z}$ from the original classifier $Q_{Y|X}$ as in equation (9).
We report results for two different scenarios in Table 2. The first is the standard SCP setting, where we assess the inefficiency of THR and APS methods, with side information $Z$ observed for 10, 30 and 100% of the instances. We redefine the classifier $f$ as in (10) to account for when $Z$ is missing, but otherwise, the SCP process remains unchanged. The second scenario is Mondrian or Group-Balanced CP [62], where one splits ${\mathcal{D}}_{cal}$ into groups and runs CP for each of them individually. In this setting, we group the calibration data points according to $Z$ and base the score function on $Q_{Y|X,Z}$. In all cases, taking the side information into account reduced the inefficiency considerably.
Table 2: Inefficiency results with side information. We report the mean prediction set size ($\pm$ std.) across 10 different calib./test splits for $\alpha=0.01$. The side information is the superclass assignment for CIFAR100 and whether the class is a digit / uppercase letter / lowercase letter for EMNIST.
@@ 7.3 Federated Learning (FL)
Table 3: Inefficiency results for conformal training in the federated setting with THR. We report the mean prediction set size ($\pm$ standard deviation) of the global federated model across 10 different calib./test splits for $\alpha=0.01$ and using THR. We use ${}_{+\text{SI}}$ to indicate the inclusion of side information. We show in bold all values within one standard deviation of the best result. Lower is better.
@@
THR THR${}_{+\text{SI}}$ THR THR${}_{+\text{SI}}$ THR THR${}_{+\text{SI}}$
CE $2.91_{\pm 0.02}$ $2.46_{\pm 0.02}$ $2.73_{\pm 0.04}$ $2.30_{\pm 0.06}$ $55.41_{\pm 1.09}$ $52.31_{\pm 1.03}$
ConfTr $4.60_{\pm 0.05}$ $3.30_{\pm 0.02}$ $10.00_{\pm 0.00}$ $10.00_{\pm 0.00}$ $\mathbf{45.60_{\pm 1.30}}$ $\mathbf{41.18_{\pm 1.16}}$
@@
Fano $\mathbf{2.63_{\pm 0.02}}$ $2.37_{\pm 0.02}$ $\mathbf{2.39_{\pm 0.07}}$ $\mathbf{2.07_{\pm 0.07}}$ $\mathbf{47.91_{\pm 1.20}}$ $\mathbf{41.19_{\pm 1.02}}$
MB Fano $2.84_{\pm 0.04}$ $2.25_{\pm 0.03}$ $\mathbf{2.52_{\pm 0.08}}$ $\mathbf{2.04_{\pm 0.07}}$ $52.94_{\pm 1.40}$ $46.97_{\pm 1.30}$
DPI $\mathbf{2.60_{\pm 0.02}}$ $2.23_{\pm 0.01}$ $2.76_{\pm 0.07}$ $2.28_{\pm 0.03}$ $52.36_{\pm 0.95}$ $48.64_{\pm 0.70}$
After convergence, we ran SCP with the final global model assuming calibration and test data sets at the server, or equivalently that the clients share their scores with the server. This reflects the best inefficiency results we can hope for with the global model, as in practice we might need to resort to privacy-preserving methods that are likely to hurt performance. See Appendix G for a discussion and extra results on other possible settings. We report inefficiency results for the global model with THR in Table 3 (see Table 14 in the appendix for APS results), where we observe similar trends to the centralized experiments in Table 1. ConfTr methods still perform well on EMNIST but struggle on CIFAR data (with the notable exception on CIFAR100, where ConfTr excelled) while our methods delivered consistent results across all data sets. This probably reflects the sensitivity of both ConfTr objectives to hyperparameters, which makes them hard to use in practice, especially in FL where hyperparameter optimization is difficult. Conversely, our bounds seem more robust to such variations, as the hyperparameters found in the centralized setting seem to translate well to the federated case.
One marked difference between Tables 1 and 3 is that the simple Fano bound (6), which lagged behind the DPI bound and its model-based counterpart in the centralized setting, achieved the best results on the federated setting. We hypothesize this could be due to overfitting of the local optimization procedures to the individual data distribution of each device, which hurts the convergence of the global model. This effect is exacerbated on CIFAR100, where we have 500 devices, each of which with very few data points. The simple Fano bound is less vulnerable to such overfitting since it relies on the main classifier to a much lesser degree. Finally, in almost all cases, our bounds outperformed the CE loss, reassuring the potential of conformal training. Moreover, the inclusion of side information reduced inefficiency in all settings, and markedly so in a few instances. This confirms the effectiveness of our side information approach in a complex and practically relevant scenario, like federated learning.